OpenStudy (anonymous):

find the y intercept and x intercept for the equation 4x + 3y = 60 whats the y intercept? whats the x intercept? type an ordered pair

OpenStudy (anonymous):

3y = 60 - 4x y = 20 - 4x/3 20 - 4x / 3 = 0 x = 15 (x intercept) y = 20 - 4(0) / 3 y = 20 (y intercept)

OpenStudy (anonymous):

ordered pair: x -> (15, 0) ordered pair: y -> (0, 20)

OpenStudy (anonymous):

NOiiicee

OpenStudy (anonymous):

fruitless the order pair above is for equation 4x + 3y = 60 if so what is the ordered pair for the one before this one 3x-6y=90

OpenStudy (anonymous):

x intercept is (30,0) y intercept is (0,-15)

OpenStudy (anonymous):

convert the equation to the following expression: x/a+y/b=1. then the a will the x- intercept and b will be the y- intercept. in this case : 4x+3y=60 4x/60+3y/60=1 x/15+y/20=1 so a=15 and b=20